# Scalable Multispecies Ion Transport in a Grid Based Surface-Electrode Trap

### v1

@@ Abstract
We present a scalable method for the control of ion crystals in a grid-based surface electrode Paul trap and characterize it in the context of transport operations that sort and reorder multispecies crystals. By combining co-wiring of control electrodes at translationally symmetric locations in each grid site with the site-wise ability to exchange the voltages applied to two special electrodes gated by a binary input, site-dependent operations are achieved using only a fixed number of analog voltage signals and a single digital input per site. In two separate experimental systems containing nominally identical grid traps, one using - crystals and the other -, we demonstrate this method by characterizing the conditional intra-site crystal reorder and the conditional exchange of ions between adjacent sites on the grid. Averaged across a multi-site region of interest, we measure sub-quanta motional excitation in the axial in-phase and out-of-phase modes of the crystals following these operations at exchange rates of 2.5 kHz. These conditional transport operations display all necessary components for sorting qubits, and could be extended to implement other conditional operations involving control fields such as gates, initialization, and measurement.
@@ I Introduction
As one of the original platforms suggested for quantum computing more than 25 years ago [1], trapped ion systems have maintained their status as a leading candidate for gate based quantum processors. Trapped ion qubits have been used to demonstrate the published world records in two-qubit gate fidelity [2, 3], single qubit gate fidelity [4] and state preparation and measurement error [5]. In conjunction with the all-to-all qubit connectivity enabled by ion transport operations [6, 7, 8], trapped ion quantum computers have raised quantum computing performance benchmarks such as quantum volume [9] to new heights [10].
Existing trapped ion quantum computers typically have achieved all-to-all connectivity using linear Paul traps, which use a combination of low bandwidth control electrodes and an rf potential to provide confinement [11, 8, 7]. In the quantum charge coupled device (QCCD) architecture, the control electrodes are leveraged to create an array of discrete confining locations (“wells”) within the linear trap. Ion pairs co-trapped in a single well are entangled through their shared motional modes [12]. Arbitrary qubit connectivity is then realized by co-trapping target qubits using operations that merge and separate wells along the trap, interleaved with swap operations that reorder qubits within a well [8, 10].
@@
A proposed solution for scaling the size of ion traps is a 2D (or 3D) grid of short linear traps connected by junctions. Proof-of-concept demonstrations of transport through such junctions have been performed [20, 21, 22, 23, 24, 25, 26, 27], including high-fidelity, low heating transport for single-ions [23] and multispecies crystals [27]. In particular, multispecies junction transport eliminates the need to separate a qubit ion from its sympathetic coolant [28], vastly simplifying the transport operations required for ion sorting [27] and mid-circuit cooling.
Without a mitigating strategy, the number of control electrodes and associated wiring elements scales linearly with the number of ions in the trap. This “wiring problem” is a common issue that is shared by many quantum computing platforms [29]. In the ion case, this scaling challenge has been partially alleviated through co-wiring, or “broadcasting”, of individual control signals across multiple surface electrodes in order to perform the same operations on multiple wells in parallel [30, 10]. This approach allows storage of more ions with fewer analog signals but does not fundamentally change the scaling, as individual control electrodes must be incorporated throughout the trapping array in order to perform arbitrary sorting. A recent proposal suggested supplementing broadcasting with a switch network between the DACs and electrodes, making the co-wiring dynamically configurable. As a result, a single, multi-bit input to the switch bank per site could determine which set of DACs is connected and therefore which sorting operation occurs [13].
In this work, we experimentally demonstrate the transport operations for sorting qubits in a grid-based rf Paul trap based on a new, scalability-focused transport primitive termed ‘center-to-left-or-right’ (C2LR). This new primitive enables independent control of the well trajectory in each grid leg using a single digital control signal per site and a fixed number of analog signals independent of grid size. The C2LR primitive can be combined with global shifts and rotations to implement higher-level operations required for a full qubit-sorting algorithm, namely site-dependent swapping of qubit pairs between adjacent legs through their shared junction and site-dependent reordering of ion crystals. In two experimental setups using nominally identical traps, we test these site-swapping operations with either - (Yb-Ba) or - (Ba-Sr) two-ion crystals and show that these crystals remain in their motional ground state at swap rates exceeding kHz, paving the way for large scale, periodic, 2D ion traps with practically realizable wiring demands.
@@ II Experimental Systems
The surface trap presented here is a 2D tiling of square cells, each made of a pair of 375 um long linear Paul segments connected at 90 degrees by an X-junction (Fig. 1). Two systems use this trap to both demonstrate the flexibility of the design but also to pursue different potential benefits: Yb-Ba is a thoroughly tested pair with existing demonstrations in junction transport [27] as well as commercial operation [8, 10], while Ba-Sr is an attractive alternative species combination requiring no ultraviolet light for trapping and control, a favorable trait for optical systems in general and integrated photonic structures in particular [31]. Individual ions are loaded through a hole through the entire trap stack using neutral flux from atomic ovens and standard photoionization techniques [32]. Crystals are formed by shifting ions away from the load hole where a combine operation joins two wells into one.
@@
The rf confinement is provided by a single, continuous electrode that uses sub-surface routing to connect otherwise separate top metal traces (Fig. 1). An rf voltage with amplitude of is applied to trap Yb-Ba crystals, while a smaller voltage of V is used to trap Ba-Sr crystals, due to the smaller total mass. In both experimental setups the rf voltage is driven at a frequency of MHz through a cryogenic rf resonator. Segmented control electrodes located on or near the trap axis provide the dynamic control over the electric potential required for transport operations. In addition, large shim electrodes, located off the trap axis, are used to compensate large-scale, static potential inhomogeneities. The shim electrodes are held at 0 V except during split/merge operations or while the ions are idling in the leg centers.
@@
The majority of the electrodes in the trap are physically co-wired according to the translational symmetry of the trap (Fig. 1). In each leg, there are 12 broadcast control electrodes, two shim electrodes and two center-to-left-or-right electrodes (C2LR, described in the next section). The junction center electrode included in this count is shared by horizontal and vertical legs. The four electrodes surrounding the junction center are independently wired for testing purposes but are counted among the broadcast electrodes as they are tied together in software for all presented operations. The C2LR electrodes are located on either side of the leg center electrode and can be controlled independently. While the trap is symmetric under 90 degree rotation, vertical and horizontal legs are wired independently.
@@ III Center-to-left-or-right primitive
At the core of our two dimensional ion transport protocol is the C2LR operation, where crystals are shifted left or right depending on how two voltages, and , are mapped onto the two C2LR electrodes, and (inset Fig. 2(a)). If the mapping between electrodes and voltage sources is fixed, as is the case in the experimental systems here, then C2LR in conjunction with broadcasting reduces the number of additional analog signals per zone required to control the trap down to two. However, a cross-over switch (double pole, single throw) in each leg could be used to dynamically control the mapping between the two voltages and C2LR electrodes with a single binary input. In this case, transport in an site trap could be controlled with a fixed number of analog signals (here, 27), digital inputs, and crossover switches.
@@
To generate the C2LR primitive, we impose mirror symmetry about the center of the legs on the applied voltages, tying pairs of electrodes on opposite sides of the center together in software. Crucially, the voltages and , applied to and respectively, are permitted to vary independently (inset Fig. 2(a)). With this single-exception mirror-symmetry constraint, we solve for a linear transport segment [27] that shifts the ions 63 µm to the right from the zone center (Fig. 2(a) and (c)). The wells are constrained such that the axial center-of-mass frequency remains constant during the transport, at MHz for Yb-Ba and MHz for Ba-Sr, while the radial mode frequencies are free to vary (Fig. 2(b)).
@@
The axial symmetry of the electrode geometry about the leg center in conjunction with the constraints described above implies that exchanging the voltages applied to the the C2LR electrodes (, ), will move the ion crystal to the left with an equivalent but mirrored trajectory compared to the as-solved rightward shift (Fig. 2(a)). In this demonstration, the re-mapping is done in software, though the outcome is identical to if a cross-over switch were used. As the horizontal and vertical legs are symmetric but independently wired, there exist equivalent operations for conditional upward or downward shifts that can be implemented in parallel with the conditional horizontal shifts described above. (Fig. 2(c)).
A method for scalable ion sorting with similar ambitions was recently proposed by Malinowski et al [13] with two key differences from the experimental demonstration here. First merge, swap, and split operations are fundamental primitives in their sorting algorithm, while the proposal here uses only linear transport through junctions. Compared to the 1D merge, swap, and split strategy, these linear shifts–even through junctions [27]–are generally faster, more reliable [33], and less sensitive to miscalibration and stray fields [34, 35]. Second, the proposal in Ref. [13] uses a switchable input on every electrode in a conditional zone, a multi-bit conditional signal, and a separate set of DACs for each condition. Our method, in its full realization, uses one binary signal and one crossover switch per zone with no additional DAC channels. Despite the increased constraints, we show that there is no sacrifice in performance; arbitrary sorting is possible with low motional excitation and high transport speeds.
@@
To demonstrate the performance of the C2LR protocol, we prepare Yb-Ba and Ba-Sr crystals in their motional ground state through a combination of Doppler, electromagnetically-induced-transparency [36] and sideband cooling. This cooling brings all motional modes to near their quantum ground state, after which we apply C2LR operations in either direction followed by its time reverse to return the ions to the zone center for measurement. We then perform thermometry using the asymmetry of the motional sidebands to measure the resulting excitation [37, 38]. The measurement ion is for Yb-Ba crystals and for Ba-Sr, and numerical analysis assumes a thermal phonon distribution with mean . We focus on the the axial in-phase (XCOM) and axial out-of-phase (XSTR) modes as they are the most relevant modes for current gating schemes [8].
@@
For all presented Yb-Ba measurements, transport sequences begin and end in one of the six horizontal legs where ion fluorescence can be collected with our specific imaging apparatus, while for all Ba-Sr measurements, transport sequences begin and end in the two central vertical legs where all desired operations can be tested while cooling and measurement are calibrated in the minimal number of locations. These two cases of measurement location, while a result of particularities of the experimental setups, nevertheless demonstrates that the trap itself is capable of high quality operation in both vertical and horizontal legs. Note that the axial modes of a crystal located in a vertical leg retain the labels XCOM and XSTR despite being rotated in the lab frame.
@@
In Fig. 2(d), we show the resulting thermal occupation of XCOM and XSTR as a function of speed averaged over all locations and both direction, with this occupation remaining low up to 8 m/s for Ba-Sr and to at least 10 m/s for Yb-Ba. The resulting occupation is dominated by how cold we can prepare the crystals ( quanta, quanta), excluding for Ba-Sr XCOM above 8 m/s. In Fig. 2(e) we show that the occupancy increases linearly as a function of the number of round trips when running the transport sequence at a fixed speed of 6 m/s, suggesting an incoherent excitation source. The values of the fits are displayed in Tab. 1.
@@
All aggregated measurements, including C2LR measurements above and swap-or-stay measurements below, calculate uncertainties on the average over the individual measurements using one standard deviation on the mean . These measurements include variation in shift direction and measurement location.
IV swap-or-stay
@@
We use the C2LR primitive as the key conditional step in a 2D sorting operation that we term ‘swap-or-stay’. The swap-or-stay sequence acts on the entire trap, swapping a target set of diagonally adjacent qubit pairs while returning the others to their initial locations after only intra-leg excursions. Swap-or-stay begins with a C2LR operation and ends with its time reverse. Between the conditional shifts are global transports that primarily transport ions into and out of the junction during which the C2LR electrodes act as generic broadcast electrodes. As these intermediate steps are global, they must be agnostic to the initial C2LR displacement direction and therefore accomplish the swap and stay operations in parallel.
@@
There are four swap-or-stay variations, each acting on a pair of adjacent crystals but across a different diagonal; for illustration, consider the swap between the crystal that is to the left (pair 1) and the one below (pair 2) the junction (Fig. 3). The swap begins with a conditional shift towards the junction center for both pairs, i.e. right for pair 1 and up for pair 2, followed by a continued upward shift sending pair 2 into the junction. Next, a rightward shift simultaneously brings pair 2 out of the junction while pair 1 moves in. After this, pair 1 shifts down towards its destination, followed by a leftward shift back through the junction by pair 2. The swap concludes with the reverse of the initial conditional shifts.
@@
In the case of Yb-Ba, there is an additional crystal axis rotation between junction transport operations. This rotation reorients the crystal in the plane of the trap so that the crystal is aligned with the subsequent linear transport direction with Yb facing the junction. With trajectories and potential well constraints chosen here combined with its larger mass ratio, the Ba-Sr crystals orient vertically at the junction center [27], thereby rendering these rotations irrelevant.
@@
Because the intermediate global transport supports either choice of initial C2LR direction, these global shifts must begin with two wells on each leg of the grid, one at each potential endpoint of the C2LR operation. The C2LR condition chooses which of the wells are occupied after the first step or brought back to the center after the penultimate one. This multi-well transport through junctions in a broadcast grid presents a challenging problem given the demands of multispecies junction transport and the limited number of electrodes per zone. Through judicious prescription of constraints on the potential wells, we successfully generated solutions for the required global shifts with constraints on the wells that enter junctions based upon previous work [27].
@@
We then characterize the motional excitation and achievable speed of the four swap-or-stay permutations throughout the region of interest. A crystal is moved sequentially to each location, where it is cooled close to its motional ground state. A doubled swap-or-stay operation at rate is then executed, which unconditionally returns the ions to the measurement zone, and the resulting excitation of the axial motional modes is characterized with sideband asymmetry. The total swap time consists of the active transport duration, interpolation time between steps (10 µs per step, Fig. 3), and, for Yb-Ba, rotation duration (20 µs, Fig. 4(e)). The varied parameter is the transport duration, determined by the average speed and the total effective distance traveled of 750 µm, twice the on-grid distance between sites. When testing the doubled swap operations, both crystal types perform the same round trip operation, differing only in where the round trip begins (Fig. 4(f)).
@@
The measured excitation across the swap-or-stays of interest (Fig. 4(a-d)) clearly demonstrates that we can perform the operation for both Yb-Ba and Ba-Sr across a range of rates and grid locations while maintaining the crystal near its motional ground state. Furthermore, the spatial uniformity of the performance illustrates the high degree of translation symmetry achieved in the trap. For Yb-Ba, we measure the fastest swap before substantial motional excitation at a rate of 2.5 kHz, corresponding to linear transport speeds of 3.3 m/s, while for Ba-Sr this occurs at 3.2 kHz and a speed of 3.0 m/s. The average resulting excitations are presented in Tab. 2. We do not subtract off the substantial contribution of imperfect XCOM ground state cooling (approximately 0.1-0.2 quanta) as the mode occupations mix considerably during junction transits [27]. This mixing does imply that measurement of the axial modes serves as an effective gauge of the radial mode excitation.
@@
Several features worthy of note emerge from these measurements (Fig. 4(a)-(d)). For large , the excitation sharply increases, a sign of the onset of coherent excitation caused by fast transport [27]. For longer transport times ( kHz), the resulting excitation is dominated by an effective heating rate that is independent of operation. In this regime, the motional excitation scales with total transport time , particularly for XCOM. This effective heating rate during the transport operations matches closely with the measured static heating rate (Fig. 6) with an offset from imperfect initial cooling. This coarse agreement demonstrates the act of transport itself contributes little motional excitation until speeds increase to the onset of coherent excitation.

### v2

# Scalable Multispecies Ion Transport in a Grid-Based Surface-Electrode Trap

@@ Abstract
Quantum processors based on linear arrays of trapped ions have achieved exceptional performance, but scaling to large qubit numbers requires realizing two-dimensional ion arrays as envisioned in the quantum charge-coupled device (QCCD) architecture. Here we present a scalable method for the control of ion crystals in a grid-based surface-electrode Paul trap and characterize it in the context of transport operations that sort and reorder multispecies crystals. By combining cowiring of control electrodes at translationally symmetric locations in each grid site with the sitewise ability to exchange the voltages applied to two special electrodes gated by a binary input, site dependent operations can be achieved using only a fixed number of analog voltage signals and a single digital input per site. In two separate experimental systems containing nominally identical grid traps, one using - crystals and the other -, we demonstrate this method by characterizing the conditional intrasite crystal reorder and the conditional exchange of ions between adjacent sites on the grid. Averaged across a multisite region of interest, we measure subquanta motional excitation in the axial in-phase and out-of-phase modes of the crystals following these operations at exchange rates of 2.5 kHz. In this initial demonstration, the logic controlling the voltage exchange occurs in software, but the applied signals mimic a proposed hardware implementation using crossover switches. These techniques can be further extended to implement other conditional operations in the QCCD architecture such as gates, initialization and measurement.
@@ I Introduction
As one of the original platforms suggested for quantum computing more than 25 years ago [1], trapped ion systems have maintained their status as a leading candidate for gate-based quantum processors. Trapped ion qubits have been used to demonstrate the published world records in two-qubit gate fidelity [2, 3], single-qubit gate fidelity [4], and state preparation and measurement error [5]. In conjunction with the all-to-all qubit connectivity enabled by ion transport operations [6, 7, 8], trapped ion quantum computers have raised quantum computing performance benchmarks such as quantum volume [9] to new heights [10].
Existing trapped ion quantum computers typically have achieved all-to-all connectivity using linear Paul traps, which use a combination of low-bandwidth control electrodes and an rf potential to provide confinement [11, 8, 7]. In the quantum charge-coupled device (QCCD) architecture, the control electrodes are leveraged to create an array of discrete confining locations (“wells”) within the linear trap. Ion pairs cotrapped in a single well are entangled through their shared motional modes [12]. Arbitrary qubit connectivity is then realized by cotrapping target qubits using operations that merge and separate wells along the trap, interleaved with swap operations that reorder qubits within a well [8, 10].
@@
A proposed solution for scaling the size of ion traps is a 2D (or 3D) grid of short linear traps connected by junctions. Proof-of-concept demonstrations of transport through such junctions have been performed [20, 21, 22, 23, 24, 25, 26, 27], including high-fidelity, low heating transport for single ions [23] and multispecies crystals [27]. In particular, multispecies junction transport eliminates the need to separate a qubit ion from its sympathetic coolant [28], vastly simplifying the transport operations required for ion sorting [27] and midcircuit cooling.
Without a mitigating strategy, the number of control electrodes and associated wiring elements scales linearly with the number of ions in the trap. This “wiring problem” is a common issue that is shared by many quantum computing platforms [29]. In the ion case, this scaling challenge has been partially alleviated through co-wiring, or “broadcasting,” of individual control signals across multiple surface electrodes in order to perform the same operations on multiple wells in parallel [30, 10]. This approach allows storage and sorting of more ions with fewer analog signals but does not fundamentally change the scaling with system size, as individual control electrodes must be incorporated throughout the trapping array in order to perform arbitrary sorting. Recent proposals have suggested [13, 31] adding an array of switches in order to dynamically change the mapping between control signals and electrodes based on binary inputs, effectively making the cowiring reconfigurable at each sorting step to achieve both scalability and sitewise control.
In this work, we experimentally demonstrate the transport operations for sorting qubits in a grid-based rf Paul trap based on a new, scalability-focused transport primitive termed ‘center to left or right’ (C2LR). This primitive enables independent control of the well trajectory in each grid leg using a single digital control signal per site and a fixed number of analog signals independent of grid size. The C2LR primitive can be combined with global shifts and rotations to implement higher-level operations required for a full qubit-sorting algorithm, namely site dependent swapping of qubit pairs between adjacent legs through their shared junction and site dependent reordering of ion crystals. In two experimental setups using nominally identical traps, we test these site-swapping operations with either - (Yb-Ba) or - (Ba-Sr) two-ion crystals and show that these crystals remain in their motional ground state at swap rates exceeding kHz, paving the way for large scale, periodic, 2D ion traps with practically realizable wiring demands.
@@ II Experimental Systems
The surface trap presented here is a 2D tiling of square cells, each made of a pair of long linear Paul segments (”legs”) connected at 90 degrees by an X-junction (Fig. 1). Two systems use this trap to both demonstrate the flexibility of the design but also to pursue different potential benefits: Yb-Ba is a thoroughly tested pair with existing demonstrations in junction transport [27] as well as commercial operation [8, 10], while Ba-Sr is an attractive alternative species combination requiring no ultraviolet light for trapping and control, a favorable trait for optical systems, in general, and integrated photonic structures in particular [32]. Individual ions are loaded through a hole through the entire trap stack using neutral flux from atomic ovens and standard photoionization techniques [33]. Crystals are formed by shifting ions away from the load hole where a merge operation joins two wells into one.
@@
The rf confinement is provided by a single, continuous electrode that uses subsurface routing to connect otherwise separate top metal traces (Fig. 1). An rf voltage with amplitude of is applied to trap Yb-Ba crystals, while a smaller voltage of V is used to trap Ba-Sr crystals, due to the smaller total mass. In both experimental setups the rf voltage is driven at a frequency of MHz through a cryogenic rf resonator. Segmented control electrodes located on or near the trap axis provide the dynamic control over the electric potential required for transport operations. In addition, large shim electrodes, located off the trap axis, are used to compensate large-scale, static potential inhomogeneities. The shim electrodes are held at 0 V except during split and merge operations or while the ions are idling in the leg centers. Shimming voltages may be necessary for sorting in larger grids or to meet gating requirements. These voltages, while necessarily site dependent, can be provided by low-bandwidth sources and multiplexed using switches and a sample-hold technique to greatly reduce their cost [13].
@@
The majority of the electrodes in the grid are physically cowired according to the translational symmetry of the trap (see the lettering scheme of electrodes in Fig. 1). In each leg, there are 12 broadcast control electrodes, two shim electrodes and two center to left or right electrodes (C2LR, described in the next section). The junction center electrode included in this count is shared by horizontal and vertical legs, but otherwise all vertical and horizontal electrodes have separate physical wiring. The electrodes directly adjacent to each junction-center electrode in Fig. 1 are independently wired for testing purposes but are counted among the broadcast electrodes as their voltages are tied together in software for all presented operations. The C2LR electrodes are located on either side of the leg center electrode and can be controlled independently.
We engineer a highly translationally symmetric rf pseudopotential in the zones highlighted in Fig. 1 by both extending the periodic geometry outward beyond the region of interest and by optimizing the shape of the terminating boundary of the rf electrode. We focus our measurements on the central six horizontal and two vertical legs (dashed line in Fig. 1) of the 20-junction grid where numerical simulations indicate the deviation in height of the rf potential minimum from the mean varies by less than 0.5%.
For all transport operations we generate a numerical electrostatic model of the entire trap top metal and use a least-squares optimization procedure to solve for a transport waveform (see Appendix B for more details) in a single unit tile of the trap, namely the tile containing legs (2,0) and (1,1). We then verify in simulation that the waveform broadcasts with sufficient performance across all unit tiles of the central portion of the grid shown in Fig. 1 before testing on the experiment.
For cooling, split or merge and measurement operations the ions are held at the center of a leg at a height of , though this height varies during transport by up to several microns.
@@ III Center-to-left-or-right primitive
At the core of our two-dimensional ion transport protocol is the C2LR operation, where crystals are shifted left or right depending on how two voltages, and , are mapped onto the two C2LR electrodes, and (inset Fig. 2(a)). If the mapping between electrodes and voltage sources is fixed, as is the case in the experimental systems here, then C2LR in conjunction with broadcasting reduces the number of additional analog signals per zone required to control the trap down to two. However, a crossover switch (double pole, single throw) in each leg could be used to dynamically control the mapping between the two voltages and C2LR electrodes with a single binary input. In this case, transport in an -site trap could be controlled with a fixed number of analog signals (here, 27), digital inputs, and crossover switches.
@@
To generate the C2LR primitive, we impose mirror symmetry about the center of the legs on the applied voltages, tying pairs of electrodes on opposite sides of the center together in software. Crucially, the voltages and , applied to and respectively, are permitted to vary independently (inset Fig. 2(a)). With this single-exception mirror-symmetry constraint, we solve for a linear transport segment [27] that shifts the ions 62.5 µm to the right of the zone center (Fig. 2(a) and (c)). Shifting by this 1/6 of the leg length maximizes the spacing between neighboring wells in subsequent transport operations such that crystals are never closer than 1/3 of the leg (), enough distance to neglect interaction between ions in separate wells (Appendix D.4) in a densely loaded trap. The wells are constrained such that the axial center-of-mass frequency remains constant during the transport, at MHz for Yb-Ba and MHz for Ba-Sr, while the radial mode frequencies are free to vary (Fig. 2(b)).
@@
The axial symmetry of the electrode geometry about the leg center in conjunction with the constraints described above implies that exchanging the voltages applied to the the C2LR electrodes (, ), moves the ion crystal to the left with an equivalent but mirrored trajectory compared to the as-solved rightward shift (Fig. 2(a)). In this demonstration, the remapping is done in software, though the outcome is identical to if a crossover switch were used. As the horizontal and vertical legs are symmetric but independently wired, there exist equivalent operations for conditional upward or downward shifts that can be implemented in parallel with the conditional horizontal shifts described above. (Fig. 2(c)).
This C2LR primitive shares the aims and means of the proposal by Malinowski et al [13], using signal broadcasting and dynamic switches to realize favorable scaling of ion transport in a grid trap. The key differences are that the Malinowski proposal is theoretical and suggests maximum flexibility is realized when every electrode has its own switch, whereas the C2LR primitive is shown experimentally to be fully capable of sorting ions in a grid trap using a single effective switch per site.
@@
To demonstrate the performance of the C2LR protocol, we prepare a single Yb-Ba or Ba-Sr crystal in its motional ground state through a combination of Doppler, electromagnetically induced transparency [34] and sideband cooling. This cooling brings all motional modes to near their quantum ground state, after which we apply C2LR operations in either direction followed by its time reverse to return the ions to the zone center for measurement. We then perform thermometry using the asymmetry of the motional sidebands to measure the resulting excitation [35, 36]. The measurement ion is for Yb-Ba crystals and for Ba-Sr, and numerical analysis assumes a thermal phonon distribution with mean . We focus on the the axial in-phase (XCOM) and axial out-of-phase (XSTR) modes as they are the most relevant modes for current gating schemes [8].
@@
For all presented Yb-Ba measurements, transport sequences begin and end in one of the six horizontal legs where ion fluorescence can be collected with our specific imaging apparatus, while for all Ba-Sr measurements, transport sequences begin and end in the two central vertical legs where all desired operations can be tested while cooling and measurement are calibrated in the minimal number of locations. These two cases of measurement location, while a result of particularities of the experimental setups, nevertheless demonstrates that the trap itself is capable of high-quality operation in both vertical and horizontal legs. Note that the axial modes of a crystal located in a vertical leg retain the labels XCOM and XSTR despite being rotated in the lab frame.
@@
In Fig. 2(d), we show the resulting thermal occupation of XCOM and XSTR as a function of speed averaged over all locations and both directions, with this occupation remaining low up to 8 m/s for Ba-Sr and to at least 10 m/s for Yb-Ba. The resulting occupation is dominated by the temperature at which we can prepare the crystals ( quanta, quanta), excluding for Ba-Sr XCOM above 8 m/s. In Fig. 2(e) we show that the occupancy increases linearly as a function of the number of round-trips when running the transport sequence at a fixed speed of 6 m/s, suggesting an incoherent excitation source. The values of the fits are displayed in Table 1.
@@
All aggregated measurements, including C2LR measurements above and swap or stay measurements below, calculate uncertainties on the average over the individual measurements using one standard deviation on the mean . These measurements include variation in shift direction and measurement location.
IV swap or stay
@@
We use the C2LR primitive as the key conditional step in a 2D sorting operation that we term ‘swap or stay’. The swap or stay sequence acts on the entire trap, swapping a target set of diagonally adjacent qubit pairs while returning the others to their initial locations after only intraleg excursions. Swap or stay begins with a C2LR operation and ends with its time reverse. Between the conditional shifts are global transports that primarily transport ions into and out of the junction during which the C2LR electrodes act as generic broadcast electrodes. As these intermediate steps are global, they must be agnostic to the initial C2LR displacement direction and, therefore, accomplish the swap and stay operations in parallel.
@@
There are four swap or stay variations, each acting on a pair of adjacent crystals but across a different diagonal; for illustration, consider the swap between the crystal that is to the left (pair 1) and the one below (pair 2) the junction (Fig. 3). The swap begins with a conditional shift toward the junction center for both pairs, i.e. right for pair 1 and up for pair 2, followed by a continued upward shift sending pair 2 into the junction. Next, a rightward shift simultaneously brings pair 2 out of the junction while pair 1 moves in. After this, pair 1 shifts down toward its destination, followed by a leftward shift back through the junction by pair 2. The swap concludes with the reverse of the initial conditional shifts.
@@
In the case of Yb-Ba, there is an additional crystal axis rotation between junction transport operations. This rotation reorients the crystal in the plane of the trap so that the crystal is aligned with the subsequent linear transport direction with Yb facing the junction. With trajectories and potential well constraints chosen here combined with its larger mass ratio, the Ba-Sr crystals orient normal to the trap plane at the junction center [27], thereby rendering these rotations irrelevant.
@@
Because the intermediate global transport supports either choice of initial C2LR direction, these global shifts must begin with two wells on each leg of the grid, one at each potential endpoint of the C2LR operation. The C2LR condition chooses which of the wells are occupied after the first step or brought back to the center after the penultimate one. This multiwell transport through junctions in a broadcast grid presents a challenging problem given the demands of multispecies junction transport and the limited number of electrodes per zone. Through judicious prescription of constraints on the potential wells, we successfully generate solutions for the required global shifts with constraints on the wells that enter junctions based upon previous work [27].
@@
We then characterize the motional excitation and achievable speed of the four swap or stay permutations throughout the region of interest. A crystal is moved sequentially to each location, where it is cooled close to its motional ground state. A doubled swap or stay operation at rate is then executed, which unconditionally returns the ions to the measurement zone, and the resulting excitation of the axial motional modes is characterized with sideband asymmetry. The total swap time consists of the active transport duration, interpolation time between steps (10 µs per step, Fig. 3), and, for Yb-Ba, rotation duration (20 µs, Fig. 4(e)). The varied parameter is the transport duration, determined by the average speed and the total effective distance traveled of 750 µm, twice the on-grid distance between sites. When testing the doubled swap operations, both crystal types perform the same round-trip operation, differing only in where the round-trip begins (Fig. 4(f)).
@@
The measured excitation across the swap or stays of interest (Fig. 4(a-d)) clearly demonstrates that we can perform the operation for both Yb-Ba and Ba-Sr across a range of rates and grid locations while maintaining the crystal near its motional ground state. Furthermore, the spatial uniformity of the performance illustrates the high degree of translation symmetry achieved in the trap. For Yb-Ba, we measure the fastest swap before substantial motional excitation at a rate of 2.5 kHz, corresponding to linear transport speeds of 3.3 m/s, while for Ba-Sr this occurs at 3.2 kHz and a speed of 3.0 m/s. The average resulting excitations are presented in Table 2. We do not subtract off the substantial contribution of imperfect XCOM ground state cooling (approximately 0.1-0.2 quanta) as the mode occupations mix considerably during junction transits [27]. This mixing does imply that measurement of the axial modes serves as an effective gauge of the radial mode excitation.
@@
Several features worthy of note emerge from these measurements (Fig. 4(a)-(d)). For large , the excitation sharply increases, a sign of the onset of coherent excitation caused by fast transport [27]. For longer transport times ( kHz), the resulting excitation is dominated by an effective heating rate that is independent of operation. In this regime, the motional excitation scales with total transport time , particularly for XCOM. This effective heating rate during the transport operations matches closely with the measured static heating rate (Fig. 6) with an offset from imperfect initial cooling. This coarse agreement demonstrates that the act of transport itself contributes little motional excitation until speeds increase to the onset of coherent excitation.